Use Laplace Transforms to solve the differential equation with initial conditions.

16. y + y=sin x, y(O) = 1.
Answer #1

We have: a if yet) hy(s) yct) < SY(3) –y(0) sin et) * (5909 – 400))+ Y) - >> (5+1) YC) -1 = => (5+1) Y3) = 1+ >> YS) = + stiLet > = + . - (5+1) (3²+1) A stl BS +C s²+1 c 1= A (5+1) + (5+1) (BS+C) St1 Atc=1 A+B=0; BtC =0 ; B =-A - A +C =0 c=A › A=C=